# {strip}[Uncaptioned image]

### v1

@@ 5. Study
RQ1 draws from D1 to evaluate the effectiveness of the gesture recognition model and whether participants are able to perform the gestures. RQ2 speaks to D2 to focus on the deterrence ability. RQ3 assess participant preferences for D3. RQs 1 and 2 are evaluated via in-lab usability tasks, and we answer RQ3 through interviews with the same participants. Our university’s IRB approved the study.
@@ 5.1. Materials and procedure
For the first task, an iPhone running the Hands-Off iOS app was placed in front of the participant. The gestures were shown in random order and participants had up to one minute to perform each gesture to have it recognized by the model once. After a gesture is detected, Hands-Off would move on to the next gesture. We recorded videos of participants while using Hands-Off and document for each gesture the seconds that it took for the gesture to be detected by the system.
The second task asked the participant to try to take a screenshot of the unlocked screen to the best of their ability. We instructed them to take screenshots by pressing two buttons on the iPhone. Naturally, it is difficult to hold a hand gesture and also take a screenshot at the same time, so we instructed the participant to use whatever method they can to “break” the system and try to take the photo. The participants received no additional hints or instructions beyond this. For each gesture, participants are given up to 3 minutes or until they choose to give up and move on to the next gesture. We again recorded videos of the participants, and for each gesture we documented the number of attempts.
@@
    Attempted screenshot but the screen became locked due to gesture not being detected when screenshot was taken.
@@
    …something personal you don’t want opened in public?
@@
The evaluation was performed in a private, mostly empty office set up for the experiment design. The participants sat at a desk and used a provided iPhone 15 (iOS 17.6) on which Hands-Off was installed. A stand attached to the case supported the phone, but the phone can also be laid flat on a table. We recorded video in two different ways: (a) a screen recording captured the exact on-screen interaction between the participant and Hands-Off; and, (b) a camera on a separate laptop captured the broader physical interaction of the participant and the device.
@@ 5.3.2. Getting around Hands-Off
Playing with the model. Participants prodded, poked, and pushed the gesture recognition model to find its working boundaries. Participants attempted different angles and camera orientations to find the limits of the model. One participant creatively attempted to “trick” the model with his index finger joints posing as a frame (without success). We also witnessed attempts to do the gesture to unlock the phone and then quickly try to screenshot. We found that the interlace gesture was a special case. With the exception of interlace, no other attempts to push the boundaries of the model resulted in successful screenshots.
Using parts of the body. With the hands engaged in performing the gesture, some participants tested alternative parts of the body to press the two side buttons on the iPhone. This included elbows, knees, and even feet. The use of feet and knees were successful because the hands could be free to perform the gesture. Some gestures, such as interlace and binoculars, did not engage all fingers (e.g., free pinkies or thumbs), which one participant was able to use to screenshot.
@@ 5.3.3. Gestures as social signals
Hand gestures as an emotional language. We asked participants to select the gesture they believe to be most appropriate under different social contexts (see Fig. 9). Responses showed clear patterns to the appropriateness of gestures in different situations.
@@ 6.4. Limitations and precautions
The computing community has a role to play and responsibility to bear to address digital harms. Problems like NCIM would not exist without decades of computing work that put cameras and mobile devices into nearly everyone’s hands, with few safeguards on their use. However, it is important to interrogate the limits of FIT—and any technical approach—in important societal problems. FIT cannot solve gender-based violence. No technology can. Problems like NCIM are technologically-mediated, but the social and societal contexts in which they exist have been around for millennia. The oppression of women and gender minorities, how the law treats these harms, how society treats survivor-victims, etc., are problems that technology cannot single-handedly “fix.” For computing communities with techno-solutionist norms, it is important to note that no technology—no matter how clever and well-intentioned—will completely solve complex societal problems. Additionally, feminism itself, like any set of values or principles, has flaws and limitations (e.g., early feminism prioritized gender while overlooking race, a pattern that was also observed with social technologies (Steele, 2021)).
We began the paper with the question: can interaction techniques be feminist? Our Hands-Off system demonstrates how they can be, but we should proceed with caution in this goal. System design seeks to abstract and concretize outcomes, but consent is a practice not an outcome. To realize a goal of feminist values in interaction techniques, we should not overfit to a design space of solving an interaction problem, but instead to the goal of augmenting an interaction process aligned with desired values.
7. Acknowledgements
We thank colleagues for their early feedback on this work. We also thank funders who supported this research.

### v3

@@ 5. Study
RQ1 draws from D1 to evaluate the effectiveness of the gesture recognition model and whether participants are able to perform the gestures. RQ2 speaks to D2 to focus on the deterrence ability. RQ3 assesses participant preferences for D3. RQs 1 and 2 are evaluated via in-lab usability tasks, and we answer RQ3 through interviews with the same participants. This research was reviewed by the University of Michigan Institutional Review Board and received an exemption as a benign behavioral intervention.
@@ 5.1. Materials and procedure
For the first task, an iPhone running the Hands-Off iOS app was placed in front of the participant. The gestures were shown in random order and participants had up to one minute to perform each gesture to have it recognized by the model once. After a gesture is detected, Hands-Off will move on to the next gesture. We recorded videos of participants while using Hands-Off and documented for each gesture the seconds that it took for the gesture to be detected by the system.
The second task asked the participant to try to take a screenshot of the unlocked screen to the best of their ability. We instructed them to take screenshots by pressing two buttons on the iPhone. Naturally, it is difficult to hold a hand gesture and also take a screenshot at the same time, so we instructed the participants to use whatever method they can to “break” the system and try to take the photo. The participants received no additional hints or instructions beyond this. For each gesture, participants are given up to 3 minutes or until they choose to give up and move on to the next gesture. We again recorded videos of the participants, and for each gesture, we documented the number of attempts.
@@
    Attempted screenshot but the screen became locked due to the gesture not being detected when a screenshot was taken.
@@
    …something personal you don’t want to be opened in public?
@@
The evaluation was performed in a private, mostly empty office set up for the experiment design. The participants sat at a desk and used a provided iPhone 15 (iOS 17.6) on which Hands-Off was installed. A stand attached to the case supports the phone, but the phone can also be laid flat on a table. We recorded video in two different ways: (a) a screen recording captured the exact on-screen interaction between the participant and Hands-Off; and, (b) a camera on a separate laptop captured the broader physical interaction of the participant and the device.
@@ 5.3.2. Getting around Hands-Off
Playing with the model. Participants prodded, poked, and pushed the gesture recognition model to find its working boundaries. Participants attempted different angles and camera orientations to find the limits of the model. One participant creatively attempted to “trick” the model with his index finger joints posing as a frame (without success). We also witnessed attempts to make the gesture to unlock the phone and then quickly try to screenshot. We found that the interlace gesture was a special case. With the exception of interlace, no other attempts to push the boundaries of the model resulted in successful screenshots.
Using parts of the body. With the hands engaged in performing the gesture, some participants tested alternative parts of the body to press the two side buttons on the iPhone. This included elbows, knees, and even feet. The use of feet and knees was successful because the hands could be free to perform the gesture. Some gestures, such as interlace and binoculars, did not engage all fingers (e.g., free pinkies or thumbs), which one participant was able to use to screenshot.
@@ 5.3.3. Gestures as social signals
Hand gestures as an emotional language. We asked participants to select the gesture they believed to be most appropriate under different social contexts (see Fig. 9). Responses showed clear patterns to the appropriateness of gestures in different situations.
@@ 6.4. Limitations and precautions
The computing community has a role to play and a responsibility to bear to address digital harms. Problems like NCIM would not exist without decades of computing work that put cameras and mobile devices into nearly everyone’s hands, with few safeguards on their use. However, it is important to interrogate the limits of FIT—and any technical approach—in important societal problems. FIT cannot solve gender-based violence. No technology can. Problems like NCIM are technologically-mediated, but the social and societal contexts in which they exist have been around for millennia. The oppression of women and gender minorities, how the law treats these harms, how society treats survivor-victims, etc., are problems that technology cannot single-handedly “fix.” For computing communities with techno-solutionist norms, it is important to note that no technology—no matter how clever and well-intentioned—will completely solve complex societal problems. Additionally, feminism itself, like any set of values or principles, has flaws and limitations (e.g., early feminism prioritized gender while overlooking race, a pattern that was also observed with social technologies (Steele, 2021)).
We began the paper with the question: can interaction techniques be feminist? Our Hands-Off system demonstrates how they can be, but we should proceed with caution in this goal. System design seeks to abstract and concretize outcomes, but consent is a practice not an outcome. To realize the goal of feminist values in interaction techniques, we should not overfit to a design space of solving an interaction problem, but instead to the goal of augmenting an interaction process aligned with desired values.
Acknowledgements.
This material is based upon work supported by the National Science Foundation under Grants 1763297 and 2311102.